The product of successor and predecessor of 999 is (A) 999000 (B) 998000 (C) 989000 (D) 1998

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The successor of a whole number is the number obtained by adding 1 to it.
The number which comes immediately before a particular number is called its predecessor.

Successor of 999 = 999 + 1 = 1000

Predecessor = 999–1 = 998
Then, product of successor and predecessor of 999 is = 1000 × 998

= 998000

Thanks